# Benchmarking digital quantum simulations and optimization above hundreds of qubits using quantum critical dynamics

### v1

@@ III.2 Results: 1D chain
The Trotter circuit for the time evolution of a spin chain requires at least two layers of $R_{ZZ}$ gates per Trotter layer, or time step. Since each $R_{ZZ}$ gate is transpiled to two hardware native two-qubit gates (ECR or CZ), the final circuit for a 1D Ising model has a two-qubit gate depth of $4\times N_{t}$ where $N_{t}$ is the number of time steps. Fig. 2 shows the density of defects scaling obtained from digitized QA on subsets of 100 qubits connected through an open line (see insets), which was chosen to minimize cumulative two-qubit gate errors along the line (cf. Appendix B). Owing to noise, the experimental results differ from the noiseless statevector results in Fig. 1c, which show a monotonic decrease in defect density with increasing $t_{\mathrm{f}}$. Hardware errors only make it possible to follow the KZ scaling and subsequent scaling regimes up to a certain threshold time indicated by the dashed horizontal line in Fig. 2. Beyond this point, the defect density rises again, as the noise accumulates with increasing circuit depth and the system decoheres (indicated by grey shading). This phenomenon has been observed in analog simulators based on superconducting qubits [27, 30], and predicted by continuous-time numerical simulations that incorporate dissipation [57]. Nevertheless, this is the first time that the existence of such an optimal working point has been observed in digital quantum simulations.
@@ IV Digitized quantum annealing for optimization
The hardware and algorithms benchmarked in Section III.3 using a homogeneous ferromagnetic model have applications in combinatorial optimization. In this scenario, the couplings $J_{ij}$ in Eq. 2 can be arbitrary in magnitude and sign. An optimization problem defined on this type of Hamiltonians may fall within the NP-complete complexity class, depending on the connectivity of the underlying graph, and are frequently employed as benchmarks for quantum optimization [72, 48]. It is therefore of interest to determine whether the results obtained from benchmarking the QKZM and identifying an optimal working point are transferable to an optimization context.
@@
QA has been little explored on digital quantum computers. Instead, the focus has been mostly on QAOA, an annealing-inspired variational ansatz and the most popular near-term algorithm for quantum optimization [49, 68, 73, 74]. However, optimizing the variational parameters in QAOA is costly in itself [75] and, once found, optimal variational parameters reduce QAOA to QA in the large-layer limit [50]. This does not mean that the optimal QAOA parameters can be directly derived from annealing schedules. Rather, the variational optimization results in optimized annealing schedules. Nonetheless, precisely because of this relationship, annealing schedules can be used to initialize QAOA parameters, thereby reducing the training cost of the variational ansatz [75, 76]. It is therefore sensible to study digitized QA in more detail. Previous experimental realizations of digitized QA featured only up to nine qubits [77]. Since digitized QA requires discretizing the time evolution, the most central question is what effect does the choice of the time step have on the success of annealing?
@@
The focus of this section is first to understand whether large algorithmic errors, i.e. large time steps, improve the performance of digitized QA. This effect has been observed in classical path-integral simulations of QA, where an unconverged, finite Trotter step is beneficial for tunneling between configurations [47], reconciling earlier expectations of quantum speed-up [78] with observations [72]. Second, we identify an optimal time step for digitized QA given a fixed number of circuit layers or time steps $N_{t}$. Fixing $N_{t}$ is meaningful, since, in reality, one has access only to a limited amount of resources such as the expendable number of gates, determining both the computational cost and, more importantly before achieving fault-tolerance, the amount of error introduced by hardware noise. At the same time, the success of QA relies on as large as possible annealing times $t_{\mathrm{f}}$, which is achieved by increasing the time step $\Delta t$ for a fixed number of time steps $N_{t}$. On the other hand, the errors of time evolution algorithms typically scale with $\Delta t$ [12]. PFs are derived on the assumption of small time steps $\Delta t\ll 1$ and algorithmic errors scale with $\mathcal{O}(\Delta t^{2})$ [55]. This naturally introduces a trade-off between QA performance and the algorithmic error due to $\Delta t$ [76].
@@
Since the solution to a classical optimization problem is a single bitstring, we are usually interested in individual measurement samples rather than expectation values. Therefore, we use Qiskit’s Sampler primitive [60] throughout this section to obtain individual measurement samples, i.e., bitstrings of measurement outcomes. Since many EM techniques apply only to expectation values, we only use REM in this section. Furthermore, we will compute the residual energy,
@@
IV.1 Dependence of the residual on the time step and spectral gap
We first study the dependence of the digitized QA result on the time step on a small system of a periodic 12-qubit chain. Three instances of disorder with couplings randomly sampled from $J_{ij}\in[-1,1]$ are chosen with varying minimum energy gaps between the ground and the first parity-preserving excited state since a smaller gap entails longer annealing times and makes finding a solution generally more difficult. Note that the Hamiltonian Eq. 3 and therefore its spectrum is a function of $s$. The minimum energy gap takes on its smallest value at the critical point. Furthermore, the first excited state of opposite parity becomes degenerate with the ground state as $s\rightarrow 1$. The spectra of the three different Hamiltonians as a function of $s$ are shown in Fig. 4a with minimum gaps ranging across one order of magnitude. Fig. 4b displays the residual energy after annealing the respective system from a on ibm_auckland 111ibm_auckland is one of IBM’s by now retired 27-qubit Falcon chips with fixed time step $\Delta t\in\{0.1,\ldots,2.0\}$ as a function of the number of time steps (depth). Here, we observe that the smallest residual energies are reached after approximately $5-10$ Trotter layers, owing to decoherence. For each $\Delta t$, the minimum residual energy achieved over all circuit depths is plotted in Fig. 4c. For all three systems, the lowest residual energy is obtained with a time step of $\Delta t>1$, specifically $1.51.0<\Delta t<$ 1.5. Noiseless statevector simulations confirm that this is not merely a consequence of hardware noise, as seen in Appendix C, yielding an optimal time step of $1.41.2<\Delta t<$ 1.4. However, in statevector simulation, the existence of a finite optimal time step is likely due to the finite range of depths chosen. Therefore, choosing an infinitesimal time step seems to be both (i) inefficient in noise-free simulations and (ii) impractical in real experiments. On the other hand, a too-large time step implies algorithmic errors.
To summarize, we observe a trade-off between realizing the largest possible annealing times while keeping the algorithmic error under control. Choosing an unconverged, large time step is indeed advantageous when using digitized QA for optimization. This is of further relevance in the context of initializing QAOA variational parameters mentioned previously. In the context of directly studying QAOA performance, Ref. [76] finds an optimal time step of $0.75$ for initializing variational parameters. Increasing $\Delta t$ even further, however, does not provide any benefits as shown by the statevector results. This observation is consequential for practical noisy settings and allows us to optimize the number of Trotter steps needed to reach a target annealing time.
@@ IV.2 Optimization of disordered heavy-hexagonal graph
[…]
[]./figures/res_en_torino.pdf
Figure 5: Hardware results of the residual energy dependence on the time step. a Residual energy averaged over $10^{5}$ samples obtained from QA on ibm_torino with fixed time steps $\Delta t\in\{0.5,\ldots,1.5\}$ as a function of circuit depth (number of time steps). b Minimum residual energy from a as a function of the time step, i.e., each point corresponds to the minimum over one curve in a.
We now study a heavy-hexagonal spin-lattice of 133 qubits, i.e. ibm_torino’s full chip. We consider just one realization of disorder with uniformly sampled couplings $J_{ij}\in[-1,1]$, and compute $E_{\mathrm{res}}$ as a function of circuit depth and time step as before. The ground state is again computed using CPLEX. This time, the system is too large to compute its instantaneous higher excited states, which is why we have no information about the size of its minimum energy gap. However, the comparison of different systems in the previous section indicates that the dependence of the optimal time step on the spectral gap is small. Fig. 5a shows the results for time steps $\Delta t\in\{0.5,\ldots 1.5\}$. Again, we observe that, for each time step, the minimum residual energy is obtained at around $5-10$ Trotter layers, or time steps, after which decoherence begins to dominate. This observation is compatible with both numerical predictions [57] and experimental results from analog simulation [30]. Furthermore, it is analogous to what was observed in the benchmarking experiment Fig. 3c. Fig. 5b reports the minimum $E_{\mathrm{res}}$ for each value of the time step. Also here, we observe a continuous decrease in $E_{\mathrm{res}}$ with increasing $\Delta t$ up to a time step of $1.0$ to $1.1$, after which $E_{\mathrm{res}}$ sharply increases. In conclusion, this indicates an unconverged, large time step of $\Delta t\gtrsim 1.0$ to be optimal for digital QA regardless of the system instance and graph connectivity.
@@ V Discussion and Outlook
We test our method on hardware native geometries and the resulting circuits contain only dense layers of $R_{ZZ}$ gates. However, this does not limit its predictive power since all two-qubit gates are eventually transpiled to the same hardware-native two-qubit gate. One can therefore directly transfer results in terms of two-qubit-gate depth to the simulation of more complicated models. Nonetheless, future research could devise more varied models exhibiting universal behavior that could be used analogously to the KZ scaling. For instance, circuits with interleaved two-qubit gates of different kinds that do not commute with each other such that they cannot be arranged in a dense layer of gates, as for example encountered in fermionic models. Furthermore, our method could be used to select best-qubit subsets. By measuring all nearest-neighbor $\sigma{}^{z}_{i}\sigma{}^{z}_{j}$ correlators on a given quantum processor, the defect densities of different qubit-subsets can be reconstructed in post-processing. For example, computing the density of defects for all qubit-subsets of size $N$ in this way would allow selecting the subset that best matches the KZ scaling. Moreover, the known finite-size scaling regimes for large annealing times could one day be used to benchmark very deep quantum circuits once they become accessible with advanced EMS techniques or, eventually, quantum error correction (QEC).
@@
Our benchmark extends beyond quantum many-body simulations to quantum optimization. We observe a clear correlation between the threshold number of Trotter steps before the onset of decoherence, identified using the KZ scaling, and the optimal depth when using digitized QA for combinatorial optimization. We demonstrate that, counter-intuitively, the time step should indeed be chosen at a value $\gtrsim 1.0$, resulting in significantly improved residual energy values after annealing. This means in practice that digitized QA is a very competitive quantum optimization framework as it (i) avoids a classical optimization loop and (ii) greatly reduces the runtime of QAOA [81], avoiding the costly iterative optimization, affected by hardware and shot noise [75, 82]. More generally, our results suggest that current hardware with simple EMS may support variational ansatze (for a similar class of Hamiltonians) comprising about ten Trotter-circuit layers. Our results seem to be robust against several instances of disorder with varying minimum spectral gaps as well as on systems of different sizes and connectivities. Nonetheless, future work should aim at exploring other settings such as solving dense graphs [73].
@@
In introducing this benchmark, we report some of the largest-scale digital quantum simulations as of today [14, 83, 84]. Compared to Ref. 14, we move closer to a physical system by incorporating Trotter error and time-dependent, general two-qubit gate angles, requiring a full gate decomposition of the $R_{ZZ}$ gate into hardware native two-qubit gates. However, we do not introduce sophisticated EMS such as probabilistic error amplification, which could be a goal of future work.
@@
Finally, we emphasize that the purpose of the paper is not to claim direct quantum advantage with the featured experiments, including the optimization demonstration. Rather, the benchmarks introduced here will be instrumental to verify future quantum simulations on sufficiently large hardware, beyond the reach of approximate classical methods.

### v2

@@ III.2 Results: 1D chain
The Trotter circuit for the time evolution of a spin chain requires at least two layers of $R_{ZZ}$ gates per Trotter layer, or time step. Since each $R_{ZZ}$ gate is transpiled to two hardware native two-qubit gates (ECR or CZ), the final circuit for a 1D Ising model has a two-qubit gate depth of $4\times N_{t}$ where $N_{t}$ is the number of time steps. Fig. 2 shows the density of defects scaling obtained from digitized QA on subsets of 100 qubits connected through an open line (see insets), which was chosen to minimize cumulative two-qubit gate errors along the line (cf. Appendix D). Owing to noise, the experimental results differ from the noiseless statevector results in Fig. 1c, which show a monotonic decrease in defect density with increasing $t_{\mathrm{f}}$. Hardware errors only make it possible to follow the KZ scaling and subsequent scaling regimes up to a certain threshold time. Beyond this point (indicated by grey shading), the defect density deviates from the $t^{-1/2}$ scaling and eventually increases, as the noise accumulates with increasing circuit depth as predicted by the noisy emulation of the algorithm.
@@ IV Digitized quantum annealing for optimization
The proposed benchmark provides a reasonable prediction of hardware and error mitigation capabilities for a set of problems broader than the non-equilibrium quantum dynamics of a ferromagnetic Ising model. Estimating how many circuit layers can be reliably executed before noise prevails is useful in many applications. For instance, variational algorithms [46, 83] for many-body spin systems feature circuits very similar to the ones considered in the previous section. Circuits with the same structure may also be used in quantum machine learning [84, 85, 86]. Finally, the entangling blocks used previously can be adapted to QAOA or digitized QA of disordered spin models.
@@
We now turn to combinatorial optimization, where the couplings $J_{ij}$ in Eq. 2 can be arbitrary in magnitude and sign. Optimization problems defined by this type of Hamiltonian may fall within the NP-complete complexity class, depending on the connectivity of the underlying graph, and are frequently employed as benchmarks for quantum optimization [87, 88]. It is therefore of interest to determine whether the results obtained from benchmarking the QKZM and identifying an optimal working point are transferable to an optimization context. This is not straightforward since the QKZM is connected with second-order phase transitions, while the bottlenecks in quantum annealing of hard optimization problems are due to first-order transitions in the glassy phase [53, 54, 23].
@@
QA has been little explored on digital quantum computers [89]. Instead, the focus has been mostly on QAOA, an annealing-inspired variational ansatz and the most popular near-term algorithm for quantum optimization [49, 79, 90, 91]. However, optimizing the variational parameters in QAOA is costly in itself [92] and, once found, optimal variational parameters reduce QAOA to QA in the large-layer limit [50]. This does not mean that the optimal QAOA parameters can be directly derived from annealing schedules. Rather, the variational optimization results in optimized annealing schedules. Nonetheless, precisely because of this relationship, annealing schedules can be used to initialize QAOA parameters, thereby reducing the training cost of the variational ansatz [92, 93]. It is therefore sensible to study digitized QA in more detail. Previous experimental realizations of digitized QA featured only up to nine qubits [94].
@@
Since digitized QA requires discretizing the time evolution, a central question is what effect does the choice of the time step have on the success of the annealing. In principle, a large time-step error could, counterintuitively, even improve the performance of digitized QA. This effect has been observed in classical path-integral simulations of QA, where an unconverged, finite Trotter step is beneficial for tunneling between configurations [52], reconciling earlier expectations of quantum speed-up [95] with observations [87]. While investigating this aspect we identify an optimal time step for digitized QA given a fixed number of circuit layers or time steps $N_{t}$. Fixing $N_{t}$ is meaningful since, in reality, one can only access a limited amount of resources such as the expendable number of gates, determining both the computational cost and, more importantly before achieving fault-tolerance, the amount of error introduced by hardware noise. The success of QA relies on as large as possible annealing times $t_{\mathrm{f}}$, which is achieved by increasing the time step $\Delta t$ for a fixed number of time steps $N_{t}$. On the other hand, the errors of time evolution algorithms typically scale with $\Delta t$ [12]. PFs are derived on the assumption of small time steps $\Delta t\ll 1$ and algorithmic errors scale with $\mathcal{O}(\Delta t^{2})$ [60]. This naturally introduces a trade-off between QA performance and the algorithmic error due to $\Delta t$ [93]. Note that, despite this trade-off, having a converged time step is no requirement here. Instead, the sole objective is to obtain the best possible solution, i.e., the lowest energy. This differs from Section III where we chose the largest possible time step that is still reasonably close to the continuum while allowing us to reach as-large-as-possible $t_{\mathrm{f}}$.
@@
Since the solution to a classical optimization problem is a single bitstring, we are usually interested in individual measurement samples rather than expectation values. Therefore, we use Qiskit’s Sampler primitive [71] throughout this section to obtain individual measurement samples, i.e., bitstrings of measurement outcomes. Many EM techniques apply only to expectation values. By contrast, error suppression methods such as DD and pulse-efficient transpilation apply to sampling. However, as discussed in Section III, DD does not yield substantial improvements on ibm_torino. Furthermore, IBM Heron processors, such as ibm_torino, are based on tunable couplers and thus do not allow for pulse-efficient transpilation which was designed for cross-resonance-based hardware [80]. To make results comparable across different devices, we therefore only use REM in this section. Furthermore, we will compute the residual energy,
@@
IV.1 Dependence of the residual energy on the time step and spectral gap
We first study the dependence of the digitized QA result on the time step on a small system of a periodic 12-qubit chain. Three instances of disorder with couplings randomly sampled from $J_{ij}\in[-1,1]$ are chosen with varying minimum energy gaps between the ground and the first parity-preserving excited state since a smaller gap entails longer annealing times and makes finding a solution generally more difficult. Note that the Hamiltonian Eq. 3 and therefore its spectrum is a function of $s$. The minimum energy gap takes on its smallest value at the critical point. Furthermore, the first excited state of opposite parity becomes degenerate with the ground state as $s\rightarrow 1$. The spectra of the three different Hamiltonians as a function of $s$ are shown in Fig. 4a with minimum gaps ranging across one order of magnitude. Fig. 4b displays the residual energy after annealing the respective system from a on ibm_auckland 111ibm_auckland is one of IBM’s by now retired 27-qubit Falcon chips with fixed time step $\Delta t\in\{0.1,\ldots,2.0\}$ as a function of the number of time steps (depth). Here, we observe that the smallest residual energies are reached after approximately $5-10$ Trotter layers, owing to decoherence. The observation of a minimum is compatible with both numerical predictions [63] and experimental results from analog simulation [31, 98]. Furthermore, it is analogous to what was observed in the benchmarking experiment Fig. 3c, where we also located a minimum defect density after roughly 5-10 Trotter layers.
For each $\Delta t$, the minimum residual energy achieved over all circuit depths is plotted in Fig. 4c. For all three systems, the lowest residual energy is obtained with a time step of $\Delta t>1$, specifically $1.51.0<\Delta t<$ 1.5. Noiseless statevector simulations confirm that this is not merely a consequence of hardware noise, as seen in Appendix E, yielding an optimal time step of $1.41.2<\Delta t<$ 1.4. However, in statevector simulation, the existence of a finite optimal time step is likely due to the finite range of depths chosen. Therefore, choosing an infinitesimal time step seems to be both (i) inefficient in noise-free simulations and (ii) impractical in real experiments. On the other hand, a too-large time step implies algorithmic errors.
To summarize, we observe a trade-off between realizing the largest possible annealing times while keeping the algorithmic error under control. Choosing an unconverged, large time step is indeed advantageous when using digitized QA for optimization. This is of further relevance when initializing QAOA variational parameters mentioned previously. In the context of directly studying QAOA performance, Ref. [93] finds an optimal time step of $0.75$ for initializing variational parameters. Increasing $\Delta t$ even further, however, does not provide any benefits as shown by the statevector results. This observation is consequential for practical noisy settings and allows us to optimize the number of Trotter steps needed to reach a target annealing time.
@@ IV.2 Optimization of disordered heavy-hexagonal graph
Figure 5: Hardware results of the residual energy dependence on the time step. a Residual energy of a heavy-hexagonal 133-qubit graph with coupling coefficients uniformly sampled from $J_{ij}\in[-1,1]$, averaged over $10^{5}$ samples obtained from QA on ibm_torino with fixed time steps $\Delta t\in\{0.5,\ldots,1.5\}$ as a function of circuit depth (number of time steps). b Minimum mean residual energy from a as a function of the time step, i.e., each point corresponds to the minimum over one curve in a.
We now study a heavy-hexagonal spin-lattice of 133 qubits, i.e. ibm_torino’s full chip. We consider just one realization of disorder with uniformly sampled couplings $J_{ij}\in[-1,1]$, and compute $E_{\mathrm{res}}$ as a function of circuit depth and time step as before. The ground state is again computed using CPLEX. This time, the system is too large to compute its instantaneous higher excited states, which is why we have no information about the size of its minimum energy gap. However, the comparison of different systems in the previous section indicates that the dependence of the optimal time step on the spectral gap is small. Fig. 5a shows the results for time steps $\Delta t\in\{0.5,\ldots 1.5\}$. Again, we observe that, for each time step, the minimum residual energy is obtained at around $5-10$ Trotter layers, or time steps, after which decoherence begins to dominate. Fig. 5b reports the minimum $E_{\mathrm{res}}$ for each value of the time step. Also here, we observe a continuous decrease in $E_{\mathrm{res}}$ with increasing $\Delta t$ up to a time step of $1.0$ to $1.1$, after which $E_{\mathrm{res}}$ sharply increases. In conclusion, this suggests an unconverged, large time step of $\Delta t\gtrsim 1.0$ to be optimal for digital QA regardless of the system instance and graph connectivity. Further studies will be required to explore the generality of this empirical finding.
IV.3 Consistency with the KZ benchmark
Here, we briefly summarize the relation of our large-scale optimization results in Section IV.2 to the results of Section III. In Section III, the 1D ferromagnetic system (i.e., uniform couplings) simulation on ibm_torino, see Fig. 2, indicates that about 5 Trotter steps can be executed reliably before the density of defects substantially deviates from the expected scaling (referring to the REM curve for consistency with the optimization experiments). Subsequently, at about 8 Trotter steps, the density of defects reaches a minimum before it increases due to the accumulated noise. Similar values of 3 and 8 Trotter steps, respectively, are obtained on the heavy-hex lattice, see Fig. 3.
The optimization experiment shares the same connectivity as these benchmarks but has random couplings and different time steps. Depending on the time step used, the minimum energy we measure (which is the disordered analog to the defect density) lies again in the range of 5-10 Trotter steps, see Fig. 5a. This finding is consistent with our results from Section III. For applications such as optimization or variational methods, the location of the minimum in $n_{\mathrm{def}}(t_{\mathrm{f}})$ is more relevant than the point where it starts deviating from the expected scaling.
This demonstrates the usefulness of our benchmarking routine in an application context. Suppose, for example, that one is interested in using QAOA rather than digitized QA. Since QAOA features the same circuit structure in a closed-loop classical parameter optimization, it is expensive to discover the optimal circuit depth empirically. Running the proposed benchmark beforehand allows one to directly select the optimal QAOA depth according to the hardware capabilities.
@@ V Discussion and Outlook
We test our method on hardware native geometries and the resulting circuits contain only dense layers of $R_{ZZ}$ gates. However, this does not limit its predictive power since all two-qubit gates are eventually transpiled to the same hardware-native two-qubit gate. One can therefore directly transfer results in terms of two-qubit-gate depth to the simulation of more complicated models. Nonetheless, future research could devise more varied models exhibiting universal behavior that could be used analogously to the KZ scaling. For instance, circuits with interleaved two-qubit gates of different kinds that do not commute with each other such that they cannot be arranged in a dense layer of gates, as for example encountered in fermionic models. Furthermore, our method could be used to select best-qubit subsets. By measuring all nearest-neighbor $\sigma{\mathstrut}^{z}_{i}\sigma{\mathstrut}^{z}_{j}$ correlators on a given quantum processor, the defect densities of different qubit-subsets can be reconstructed in post-processing. For example, computing the density of defects for all qubit-subsets of size $N$ in this way would allow selecting the subset that best matches the KZ scaling. Moreover, the known finite-size scaling regimes for large annealing times could one day be used to benchmark very deep quantum circuits once they become accessible with advanced EMS techniques or, eventually, quantum error correction (QEC).
@@
Our benchmark extends beyond quantum many-body simulations to quantum optimization and variational algorithms. We observe consistency between the threshold number of Trotter steps before the noise becomes prevalent, identified using the KZ scaling, and the optimal depth of digitized QA for combinatorial optimization. We demonstrate that, counter-intuitively, the time step should indeed be chosen at a value $\gtrsim 1.0$, resulting in significantly improved residual energy values after annealing. This means in practice that digitized QA could be a competitive quantum optimization framework as it (i) avoids a classical optimization loop and (ii) greatly reduces the runtime of QAOA [99], avoiding the costly iterative optimization, affected by hardware and shot noise [92, 100]. More generally, our results suggest that current hardware with simple EMS may support variational ansatze (for a similar class of Hamiltonians) comprising about ten Trotter-circuit layers. Our results seem to be robust against several instances of disorder with varying minimum spectral gaps as well as on systems of different sizes and connectivities. Nonetheless, future work should aim at exploring other settings such as solving dense graphs [90].
@@
In introducing this benchmark, we report some of the largest-scale digital quantum simulations as of today [14, 101, 102, 103]. Compared to Ref. 14, we move closer to a physical system by incorporating Trotter error and time-dependent, general two-qubit gate angles, requiring a full gate decomposition of the $R_{ZZ}$ gate into hardware native two-qubit gates. However, we do not introduce sophisticated EMS such as probabilistic error amplification, which could be a goal of future work.
@@
Finally, we emphasize that the purpose of this paper is not to claim direct quantum advantage with the featured experiments, including the optimization demonstration. For one, the sparse-graph optimization problem studied here is not hard. Other quantum or classical platforms may demonstrate better performance than we do [104]. Moreover, we do not claim that the digitized quantum annealing method is the best-performing digital algorithm. Indeed, warm-start methods [105, 106] and digitized counterdiabatic approaches [107] all improve the performance of quantum approximate optimization. Instead, the value of our optimization experiments lies in understanding its correlation with the proposed benchmark. Furthermore, the benchmarks introduced here will be instrumental to verify future quantum simulations on sufficiently large hardware, beyond the reach of approximate classical methods.